# Ludax: A GPU-Accelerated Domain Specific Language for Board Games

### v1

@@ Abstract
Games have long been used as benchmarks and testing environments for research in artificial intelligence. A key step in supporting this research was the development of game description languages: frameworks that compile domain-specific code into playable and simulatable game environments, allowing researchers to generalize their algorithms and approaches across multiple games without having to manually implement each one. More recently, progress in reinforcement learning (RL) has been largely driven by advances in hardware acceleration. Libraries like JAX allow practitioners to take full advantage of cutting-edge computing hardware, often speeding up training and testing by orders of magnitude. Here, we present a synthesis of these strands of research: a domain-specific language for board games which automatically compiles into hardware-accelerated code. Our framework, Ludax, combines the generality of game description languages with the speed of modern parallel processing hardware and is designed to fit neatly into existing deep learning pipelines. We envision Ludax as a tool to help accelerate games research generally, from RL to cognitive science, by enabling rapid simulation and providing a flexible representation scheme. We present a detailed breakdown of Ludax’s description language and technical notes on the compilation process, along with speed benchmarking and a demonstration of training RL agents. The Ludax framework, along with implementations of existing board games, is open-source and freely available.
@@ 1 Introduction
For the past 75 years, games have served as vital tests and benchmarks for artificial intelligence research. While many specific games have been completely solved schaeffer2007checkers or optimized beyond the abilities of the strongest human players campbell2002deep ; silver2017mastering , the general space of games remains a fertile ground for measuring improvements in reasoning, planning, and strategic thinking. A critical part of this progress, however, is the ability to test approaches and algorithms on a set of environments that are both diverse and computationally efficient.
@@
To help drive further games and learning research, we introduce Ludax: a domain-specific language for board games that compiles into GPU-accelerated code written in the JAX library jax2018github . Ludax draws on two main inspirations: (1) Ludii Piette_2020_Ludii , a general purpose description language for board games capable of representing more than 1400 games from throughout history and around the world, and (2) PGX koyamada2023pgx , a collection of optimized JAX-native implementations of classic board games and video games designed to facilitate rapid training and evaluation of modern reinforcement learning (RL) agents. Ludax presents a flexible and general-purpose game representation format that can be leveraged for efficient simulation and learning on modern computing hardware.
@@
Ludax currently supports two-player, perfect-information, turn-based board games played by placing pieces onto empty board cells. While this is a narrow class of games when compared to the full breadth of human designs and the set of mechanics implemented by Ludii, it is still broad enough to capture a wide range of existing games (e.g. Connect Four, Pente, Hex, …) as well as many unexplored novel games and variants that fall within that class. Further, Ludax is designed to be easily expandable – like with Ludii, implementing new game mechanics in Ludax only requires implementing new atomic components in the underlying description language. These components can then be combined compositionally with existing elements of the language to produce an entirely new range of possible games, instead of each game needing to be implemented separately.
@@
Another design goal for Ludax is ease of use, both in terms of game design and experimentation. The syntax of the description language is “ludemic” Piette_2020_Ludii – splitting game rules into clear sections governing the game’s setup, play mechanics, and end conditions. Like with Ludii, game programs in Ludax resemble English descriptions of rules (see Figure 1). Further, by leveraging the structure of the existing PGX library, environments instantiated in Ludax can be easily combined with existing frameworks for GPU-accelerated search, reinforcement learning, or evolution deepmind2020jax ; tang2022evojax . Ludax also supports a basic web interface for interactive debugging, with the aim to provide support for online interactive experiments in the future.
@@
To our knowledge, Ludax is the first game description language which compiles into GPU-accelerated code. In the following sections, we provide a detailed description of the language syntax, compilation process, and Ludax’s expressive range. We also provide speed benchmarking compared to both Ludii and PGX, as well as an initial demonstration of training learned agents. Finally, we conclude with a discussion of potential use cases and future directions. Ludax is open-source under the Apache 2.0 license, and the code is available at: https://github.com/gdrtodd/ludax.
Figure 1: Example game descriptions for Reversi, Pente, and Yavalax in the Ludax game description language, along with literal translations into natural language.
@@ 2 Related Work
While our core contribution of combining game description languages (GDLs) with hardware accelerated learning environments is novel, there is extensive research in each of these domains separately.
Game Description Languages: Game description languages have been used for many years and in a variety of domains. The Stanford GDL Love_2008_GDL ; Genesereth_2014_GGP ; Schiffel_2014_Representing ; Thielscher_2017_GDLIII is among the most influential, helping to popularize research in general game playing Pitrat_1968_GGP through its use in the International General Game Playing Competition Genesereth_2005_GGP ; Genesereth_2013_GGP . Other notable examples include VGDL ebner2013towards ; Schaul_2013_VGDL ; Schaul_2014_Extensible (primarily known from its use in the General Video Game AI framework Perez_2019_GVGAI ), RBG Kowalski_2019_Regular , Ludi Browne_2009_PhD , and its successor Ludii Piette_2020_Ludii . GDLs have also been used to describe the rules of card games font2013card as well as to represent human goals in naturalistic simulated environments davidson2022creativity ; davidson2025goals . Modern game description languages have tended to move away from a basis in formal logic in favor of greater human usability, though there are benefits in efficiency gained by the use of regular languages Kowalski_2020_Efficient .
GPU-Accelerated Environments: Recent years have seen a proliferation of learning environments implemented in the JAX library or other frameworks that enable hardware (typically GPU) acceleration. Examples include single-agent and multi-agent physics simulators freeman1brax ; Makoviychuk_2021_Isaac ; Bettini_2022_VMAS , ports of both classic and recent reinforcement learning tasks Dalton_2020_Accelerating ; gymnax2022github ; koyamada2023pgx ; matthews2024craftax , combinatorial optimization problems bonnet2023jumanji , multi-agent coordination problems rutherford2023jaxmarl , and driving simulators Gulino_2023_Waymax ; kazemkhani2025gpudrive . While these efforts have spurred significant progress and span a wide range of domains and task formulations, each of them implement a fixed environment or set of environments. As such, they cannot easily be extended to novel environments without first writing new hardware-accelerated code.
@@ 3 Description Language Details
Ludax’s game description language draws heavily on the Ludii description language, particularly in its use of “ludemic” syntax that represents game rules in terms of high-level and easily-understandable components Piette_2020_Ludii . The complete grammar file and syntax details are available in the Supplemental Material.
3.1 Equipment
The equipment section contains information about the physical components used by the game. Currently, this only specifies the size and shape of the board (i.e. whether it is square, rectangular, hexagonal, or hexagonal-rectangular). The dimensions and shape of the board are used during compilation to help pre-compute certain game-relevant properties, such as the board indices corresponding to lines of specific lengths. In future versions of Ludax, the equipment section will also detail the different pieces used by each player if the game specifies more than one.
3.2 Start Rules
The start section is an optional section that contains the rules for the game’s setup. For most games, play begins on an empty board and the start section is omitted. In some games, such as Reversi (see Figure 1, left panel), pieces are placed in a particular arrangement at the start of play.
@@ 3.3 Play Rules
Typically, the play rules of each game are the most involved, as they detail the core mechanics and dynamics of the game. The play section is itself broken into one or more subsections called “play phases.” Each phase has its own rules for player actions and turn-taking, as well as specific conditions for when to transition to another phase. Most games have only a single phase in which players alternate turns until the game is over, specified with the repeat keyword. Some games include a once-through phase that progresses through the turn order a single time before advancing to the next phase. The sequence of player turns is specified independently for each phase. For instance, Yavalax (Figure 1, right panel) begins with the first player making a single move (i.e. (once-through (P1) ...) ) before both players alternate taking two turns for the rest of the game (i.e. (repeat (P2 P2 P1 P1) ...) ).
@@
The core of each phase is a “play mechanic” that encodes the ways that players take their turns. In the context of reinforcement learning, a play mechanic specifies both the action space ($\mathcal{A}$) and the transition function ($\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S})$. At a lower level, each play mechanic also defines a “legal action mask function” that returns whether each action is valid from the current game state. Currently, Ludax supports only one kind of play mechanic: place. A place mechanic’s primary argument is a destination constraint which specifies where a piece may be placed on a given player’s turn. For many games, such as Tic-Tac-Toe, this is simply the set of empty board positions. For some games, however, the destination constraint is more involved: in Connect Four, legal actions are empty spaces that are on the bottom edge of the board or immediately above an occupied position (see subsection 3.5 for a discussion of how actions are represented more generally in Ludax). Even further, some games have what we call result constraints which require that a legal action results or doesn’t result in a board with specific properties. Yavalax and Reversi both use result constraints: the former forbids players from placing a piece that forms a line of five or that forms only a single line of four, whereas the latter requires players to place a piece in a way that “sandwiches” one or more of their opponent’s pieces in a line. Finally, a play mechanic may optionally specify one or more effects that modify the game state after the action is performed. Effects are used to handle mechanics like capturing or flipping pieces, as well as updating each player’s score (if the game uses score). Both Reversi and Pente use play effects to handle flipping and capturing pieces, respectively, with Pente also using the score as an alternate winning condition.
Throughout this section, we have been referring to various properties of a game state and relationships between pieces / positions (e.g. whether pieces are “sandwiched,” whether a line is formed, whether a piece is adjacent to another, …). These are the lowest-level component’s of Ludax’s description language and are referred to collectively as masks, functions, and predicates. A mask takes in the current game state and returns a boolean array over each position on the board. Some masks, like occupied or edge, take additional grammatical arguments which might specify a particular player or region of the board. [^1] The adjacent mask is a special case – it takes another mask as an additional argument and returns the board positions adjacent to any of the active positions in the original mask. A function similarly takes in the current game state and returns a single non-negative integer. In Ludax’s current form, line is probably the most commonly-used function – it returns the number of contiguous lines of a given player’s pieces on the board, with a specified length and orientation. Lastly, a predicate maps from a game state to a single boolean truth value. Many predicates operate over the outputs of masks and functions, such as exists or equals, though some like mover_is are computed directly from game states. Crucially, the outputs of masks, functions, and predicates can be combined compositionally using first-order logic (excluding quantification) to form more complicated expressions. So, the condition “if Player 2 makes a line of 4 in a row or a diagonal line of 3…” would be rendered as follows:
(and (mover_is P2) (or (line 4) (line 3 orientation:diagonal)))
Note that, for ease of use, Ludax automatically interprets the presence of a bare function inside a boolean operator as indicating a non-zero value. So, (line 4) is equivalent to (>= (line 4) 1).
@@ 3.4 End Rules
The last section of a game description in Ludax details the criteria that terminate a game. The end section contains one or more “end conditions” – these are applied in order, with the first condition to activate determining the ending behavior (i.e. which player wins or if the game ends in a draw). If none of the conditions activate, then the game continues. For instance, Tic-Tac-Toe includes both the end conditions (if (line 3) (mover win)) and (if (full_board) (draw)), with the draw condition only triggering if the “three in a row condition” is not met. End conditions also frequently refer to a player’s score, which is updated or set as a result of an action’s effects (see above).
@@ 3.5 Design Considerations
While Ludax draws heavily from the Ludii description language, there are some important differences which go beyond just changes in syntax. The first of these relates to how both systems represent a game’s action space. One of the design goals of Ludii is that game descriptions should resemble as much as possible the rules in natural language. In Connect Four, for instance, players take a move by dropping a piece into one of seven columns of the board, at which point the piece falls until it reaches the bottom or rests on another piece. Accordingly, the canonical representation of Connect Four in Ludii features pieces that “Drop” into the “LastColumn” chosen by the player (PGX implicitly represents the game in a similar way). As mentioned above, however, Ludax represents the action space differently: players simply place a piece onto an empty board cell, with actions that are not directly above an existing piece or the bottom of the board marked as illegal. Mechanically, the two implementations of Connect Four are identical – the difference lies in how they are encoded (especially to simulated players or reinforcement learning agents). The “column-based” representation has many advantages (it matches the physical properties of the game in real life and lowers the branching factor), but it is also game-specific. While Ludax also strives to represent game descriptions intuitively, we primarily aim to provide a unified representation format across games, such that general game-playing agents can more easily transfer knowledge and expertise from one game to another. As such, the size and form of the action space for any place-based game is determined only by the size and shape of the board. This choice is also partially motivated by the specifics of working with the JAX library (see Section 4) and has implications for benchmarking and downstream use-cases (see Section 6).
@@ 4 Compiling Game Descriptions into Game Environments
In this section, we describe the high-level approach used to map from programs in the Ludax game description language to hardware-accelerated simulation environments. While Ludax specifically instantiates board game environments using the Lark Python library, the general approach is flexible enough to be used with different domains and parsing toolkits. Broadly speaking, Ludax operates by defining the leaves of the grammatical parse tree (i.e. individual masks, functions, and predicates) as atomic functions written in JAX, which are then dynamically composed from the bottom-up to form higher-level operators used by the environment class. Consider again the following game expression:
(and (mover_is P2) (or (line 4) (line 3 orientation:diagonal)))
During compilation, the leaf-level nodes (i.e. (mover_is P2) and (line 4)) are converted into JAX functions which map from the current game state to (in this case) a boolean truth value, and those functions are then passed up the parse tree. Higher-level nodes, such as (and ...), receive the JAX functions corresponding to each of their children and return a new JAX function that also takes the game state as input and implements the appropriate operation (in this case, boolean conjunction). In pseudocode, using the Lark library’s Transformer paradigm, this looks like the following:
@@
In actuality, both the “children functions” and the combined “predicate function” must be written to be compatible with JAX’s vectorization scheme and just-in-time (JIT) compilation. This imposes a number of implementational constraints, most notably that the size and shape of all arrays must be fixed at compile time. This means, for instance, that the dimensions of the “legal action mask” (and, hence, the size of the action space in general) cannot change as the game progresses. In addition, values like the number of iterations in a loop or the positions of a lookup mask must essentially be “pre-specified.” Crucially, however, values that are determined during parsing (such as the number of children for a given node, or the value of any arguments) can be safely passed into compiled JAX functions as static constants. This fact is what allows Ludax to create JAX functions dynamically that nonetheless obey the constraints of vectorization and JIT compilation. At the top of the parse tree, these composed JAX functions are ultimately used to define the behaviors that appear in the environment’s step function, such as applying the player’s action to the board and handling move effects.
We next discuss some of the specific optimizations used by Ludax. In general, these are not global optimizations: they apply only to certain compositions of game rules and mechanics. Our approach is to deploy these optimizations when they are available and to “fall back” on slower but more general solutions when they are not.
@@
Precomputation: An important optimization used by the PGX library (and JAX environments more generally) is to express functions as batched matrix operations rather than iterative procedures. For instance, rather than checking for a line of pieces in Tic-Tac-Toe by starting at the position of the last move and scanning out in each direction (as Ludii’s implementation does), PGX hard-codes the set of board indices that correspond to each possible line of three in the game (i.e. [[0, 1, 2], [0, 3, 6], [0, 4, 7], ...]) and performs a single multi-dimensional index into the board array – if any of the of the board index triples all correspond to positions occupied by a single player, then the game is over. Ludax adopts and generalizes this approach: during parsing of line, for example, the line indices are computed with respect to the size and shape of the game board (i.e. rectangular, hexagonal, …) as well as the length and orientation of the desired line (i.e. diagonal, vertical, …). Again, because these values depend only on attributes that are determined during parsing, they can be passed into JAX functions as constants. Precomputation naturally causes a trade-off between compile-time and run-time efficiency. In our case, we opt to use precomputation whenever possible, though some masks and functions cannot be expressed this way.
@@
Dynamic State Attributes: Different games require tracking different kinds of information about the current game state. Most obviously, some games track a score for each player while others do not. When Ludax compiles a game, it automatically extracts the attributes required to instantiate a game state and omits the others, thereby reducing the memory footprint of the entire state object. More importantly, Ludax also automatically adds intermediary computations to each call of the environment’s step function that help speed up later mask, function, or predicate evaluations. For example, in Hex, the game ends when one player manages to connect two opposite sides of the board with a continuous path of their pieces. Naively, checking whether the edges of the board are linked requires the expensive step of computing the board’s connected components after each move. However, updating the board’s connected components as a result of placing a single piece can be done very efficiently (a technique used well in the PGX implementation). At compile time, Ludax determines whether a game makes use of a “connection” rule and modifies the step function to iteratively update and track the board’s connected components if so, greatly speeding up later checks. In future extensions, this functionality will be used to accommodate games with atypical or computationally expensive rules without affecting the runtime of existing games.
@@
Figure 2: Average throughput (moves per second) on various exemplar games for Ludax, Ludii, and PGX. The top row of games are implemented in all three frameworks, while the bottom row of games are implemented only in Ludax and Ludii. Speeds for Ludax and PGX are reported for 500 episodes of various batch sizes on a workstation with a single NVIDIA 4090 GPU and 32 CPU cores, while speeds for Ludii are reported for parallel execution on the same workstation across 1, 16, and 32 threads. Error bars are standard deviations calculated over the 500 episodes.

### v2

# Ludax: A GPU-Accelerated Description Language for Board Games

@@ Abstract
Games have long been used as benchmarks and testing environments for research in artificial intelligence. A key step in supporting this research was the development of game description languages: frameworks that compile domain-specific code into playable and simulatable game environments, allowing researchers to generalize their algorithms and approaches across multiple games without having to manually implement each one. More recently, progress in reinforcement learning (RL) has been greatly aided by advances in hardware acceleration. Libraries like JAX allow practitioners to take full advantage of cutting-edge computing hardware, often speeding up training and testing by orders of magnitude. Here, we present a synthesis of these strands of research: a domain-specific language for board games which automatically compiles into hardware-accelerated code. Our framework, Ludax, combines the generality of game description languages with the speed of modern parallel processing hardware and is designed to fit neatly into existing deep learning pipelines. We envision Ludax as a tool to help accelerate games research generally, from RL to cognitive science, by enabling rapid simulation and providing a flexible representation scheme. We present a detailed breakdown of Ludax’s description language and technical notes on the compilation process, along with speed benchmarking and a demonstration of training RL agents.
@@ 1 Introduction
For the past 75 years, games have served as vital tests and benchmarks for artificial intelligence research. While many specific games have been completely solved (schaeffer2007checkers) or optimized beyond the abilities of the strongest human players (campbell2002deep; silver2017mastering), the general space of games remains a fertile ground for measuring improvements in reasoning, planning, and strategic thinking. A critical part of this progress, however, is the ability to test approaches and algorithms on a set of environments that are both diverse and computationally efficient.
@@
To help drive further games and learning research, we introduce Ludax: a domain-specific language for board games that compiles into GPU-accelerated code written in the JAX library (jax2018github). Ludax draws on two main inspirations: (1) Ludii (Piette_2020_Ludii), a general purpose description language for board games already used to represent more than 1400 games from throughout history and around the world, and (2) PGX (koyamada2023pgx), a collection of optimized JAX-native implementations of classic board games and video games designed to facilitate rapid training and evaluation of modern reinforcement learning (RL) agents. Ludax presents a flexible and general-purpose game representation format that can be leveraged for efficient simulation and learning on modern computing hardware.
@@
Ludax currently supports two-player, perfect-information, turn-based board games played by placing, capturing, and moving pieces. This set of mechanics is broad enough to capture a wide range of existing games (e.g. Connect Four, Pente, Hex, …) as well as many unexplored novel games and variants that fall within that class. Further, Ludax is designed to be easily expandable – like with Ludii, implementing new game mechanics in Ludax only requires implementing new atomic components in the underlying description language. These components can then be combined compositionally with existing elements of the language to produce an entirely new range of possible games, instead of each game needing to be implemented separately.
@@
Another design goal for Ludax is ease of use, both in terms of game design and experimentation. The syntax of the description language is “ludemic” (Piette_2020_Ludii) – splitting game rules into clear sections governing the game’s setup, play mechanics, and end conditions. Like with Ludii, game programs in Ludax resemble English descriptions of rules (see Figure 1). Further, by leveraging the structure of the existing PGX library, environments instantiated in Ludax can be easily combined with existing frameworks for GPU-accelerated search, reinforcement learning, or evolution (deepmind2020jax; tang2022evojax). Ludax also supports a basic web interface for interactive debugging and potential user-studies.
Ludax is fundamentally a platform for accelerating board game research. In an era of increasingly complicated tasks and benchmarks, relatively simple board games may seem to be less interesting research domains (especially as many games have been more-or-less “solved” by modern methods). However, Ludax is not just a collection of new tasks. By decoupling rapid execution from the intensive process of writing new environment code, Ludax can power new research in a variety of directions. For instance, Ludax can be used to analyze RL generalization (soemers2025environment) by defining a wide range of modifications for a target game task (akin to a platform like Minihack (Samvelyan_2021_MiniHack)) or help improve studies of game generation by enabling the rapid evaluation of procedurally-generated rulesets (todd2024gavel; collins2025generation). Finally, Ludax can help advance recent research into world modeling (ying2025assessing) by (1) providing a wide and easily-refreshable set of environments to test on efficiently and (2) allowing automated systems to propose and refine world models in these “novel games” by writing high-level and semantically-meaningful DSL code.
@@
To our knowledge, Ludax is the first board game description language which compiles into GPU-accelerated code. In the following sections, we provide a detailed description of the language syntax, compilation process, and Ludax’s expressive range. We also provide speed benchmarking compared to both Ludii and PGX, as well as an initial demonstration of training learned agents. Finally, we conclude with a discussion of potential use cases and future directions. Ludax is open-source under the Apache 2.0 license, and the code is available at: https://github.com/gdrtodd/ludax.
Figure 1: Natural language description of Reversi along with its corresponding translation into Ludax. Ludax uses “ludemic” syntax that represents high-level game components as separate program sections and aims to be easily interpretable to non-experts.
@@ 2 Related Work
Game Description Languages: Game description languages have been used for many years and in a variety of domains. The Stanford GDL (Love_2008_GDL; Genesereth_2014_GGP; Schiffel_2014_Representing; Thielscher_2017_GDLIII) is among the most influential, helping to popularize research in general game playing (Pitrat_1968_GGP) through its use in the International General Game Playing Competition (Genesereth_2005_GGP; Genesereth_2013_GGP). Other notable examples include VGDL (ebner2013towards; Schaul_2013_VGDL; Schaul_2014_Extensible) (primarily known from its use in the General Video Game AI framework (Perez_2019_GVGAI)), RBG (Kowalski_2019_Regular), Ludi (Browne_2009_PhD), and its successor Ludii (Piette_2020_Ludii). GDLs have also been used to describe the rules of card games (font2013card) as well as to represent human goals in naturalistic simulated environments (davidson2022creativity; davidson2025goals). Modern game description languages have tended to move away from a basis in formal logic in favor of greater human usability, though there are benefits in efficiency gained by the use of regular languages (Kowalski_2020_Efficient).
GPU-Accelerated Environments: Recent years have seen a proliferation of learning environments implemented in the JAX library or other frameworks that enable hardware (typically GPU) acceleration. Examples include single-agent and multi-agent physics simulators (freeman1brax; Makoviychuk_2021_Isaac; Bettini_2022_VMAS), ports of both classic and recent reinforcement learning tasks (Dalton_2020_Accelerating; gymnax2022github; koyamada2023pgx; matthews2024craftax), combinatorial optimization problems (bonnet2023jumanji), multi-agent coordination problems (rutherford2023jaxmarl), and driving simulators (Gulino_2023_Waymax; kazemkhani2025gpudrive). While these efforts have spurred significant progress and span a wide range of domains and task formulations, each of them implement a fixed environment or set of environments. As such, they cannot easily be extended to novel environments without first writing new hardware-accelerated code. Ludax stands alongside a number of description languages for other domains (e.g. probabilistic programming, planning, single-player puzzles) that leverage JAX for efficient execution (chandra2025memo; gimelfarb2024jaxplan; earle2025autoverse; earle2025puzzlejaxbenchmarkreasoninglearning).
@@ 3 Description Language Details
Ludax’s game description language draws heavily on the Ludii description language, particularly in its use of “ludemic” syntax that represents game rules in terms of high-level and easily-understandable components (Piette_2020_Ludii). The complete grammar file and syntax details are available in the Supplemental Material.
3.1 Equipment and Start Rules
The equipment section contains information about the physical components used by the game. This includes the size and shape of the board (e.g. whether it is rectangular or hexagonal), the piece types used by players during the game, and optionally any “board regions” that are referred to by later rules (e.g. particular tiles that are blocked off from play). The start section is an optional section that contains the rules for the game’s setup. For most games, play begins on an empty board and the start section is omitted. In some games, such as Reversi (see Figure 1), pieces are placed in a particular arrangement at the start of play.
@@ 3.2 Play Rules
Typically, the play rules of each game are the most involved, as they detail the core mechanics and dynamics of the game. The play section is itself broken into one or more subsections called “play phases.” Each phase has its own rules for player actions and turn-taking, as well as specific conditions for when to transition to another phase. Most games have only a single phase in which players alternate turns until the game is over, though Ludax also supports games with more complicated dynamics.
@@
The core of each phase is a “play mechanic” that encodes the ways that players take their turns. In the context of reinforcement learning, a play mechanic specifies both the action space ($\mathcal{A}$) and the transition function ($\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S})$. At a lower level, each play mechanic also defines a “legal action mask function” that returns whether each action is valid from the current game state. Ludax supports two kinds of play mechanics: place and move, the latter of which is broken down into step, hop, and slide. Each play mechanic takes arguments that define exactly how it is implemented. For instance, a place rule takes a destination constraint which can be used to restrict where pieces can be placed, as in Connect Four. Arguments to hop can be used to specify which pieces can be hopped over and in which directions, and the different movement types can be composed to create intricate rulesets like English Draughts (see Appendix A). Finally, a play mechanic might optionally define one or more effects that modify the game state after the action is performed. Effects are used to handle mechanics like capturing or flipping pieces, as well as updating each player’s score if the game uses it – see Reversi again for an example.
High-level operators like place and slide eventually ground out to low-level components that encode individual game mechanics. These components, referred to as masks, functions, and predicates based on their return values, take in the current game state and compute a single, specific game property. For instance, the commonly-used line function returns the number of contiguous lines of a particular piece of a specified length and orientation while the custodial mask checks for the presence of a “custodial” arrangement in which a line of pieces belonging to one player are “sandwiched” by a pair of pieces belonging to the other player. Ludax implements a wide range of low-level mechanics, and crucially these mechanics can be combined compositionally using first-order logic (excluding quantification) to form more complicated expressions. As an example, the condition “if a player makes a line of 4 stones in a row or a 2 by 2 square of stones…” would be rendered as follows:
(or (line ‘‘stone’’ 4) (pattern ‘‘stone’’ (square 2)))
@@ 3.3 End Rules
The last section of a game description in Ludax details the criteria that terminate a game. The end section contains one or more “end conditions” which award a victory or loss to one to more players based on a binary predicate (e.g. forming a line or reaching a score threshold). A game can have multiple end conditions, which are then checked in order (so that win/loss conditions take precedence over draw checks, for instance).
@@ 3.4 Design Considerations
While Ludax draws heavily from the Ludii description language, there are some important differences which go beyond just changes in syntax. Take Connect Four as an example: Ludii aims to represent the game mechanically in a way that matches the natural language description of its rules as closely as possible. Accordingly, the canonical representation of Connect Four in Ludii features pieces that “Drop” into the “LastColumn” chosen by the player. Ludax represents the action space differently – players simply place a piece onto an empty board cell, with actions that are not directly above an existing piece or the bottom of the board marked as illegal. Mechanically, the two implementations of Connect Four are identical – the difference lies in how they are encoded (especially to simulated players or reinforcement learning agents). While Ludax, like Ludii, strives to represent game descriptions intuitively, we primarily aim to provide a unified representation format across games so that general game-playing agents can more easily transfer knowledge and expertise from one game to another. As such, the size and shape of the action space for any game in Ludax is determined exclusively by the size of the board and the high-level play mechanics, regardless of how those mechanics might be translated into natural language. This choice is also partially motivated by the specifics of working with the JAX library (see Section˜4) and has implications for benchmarking and downstream use-cases (see Section˜6).
@@ 4 Compiling Game Descriptions into Game Environments
Ludax uses the Lark Python library to parse game descriptions into JAX environments, but the general approach is flexible enough to be used with different domains and parsing toolkits. Broadly speaking, Ludax operates by defining the leaves of the grammatical parse tree (i.e. individual masks, functions, and predicates) as atomic functions written in JAX, which are then dynamically composed from the bottom-up to form higher-level operators used by the environment class. Consider again the following game expression:
(or (line ‘‘stone’’ 4) (pattern ‘‘stone’’ (square 2)))
During compilation, the leaf-level nodes (line ...) and (pattern ...) are converted into JAX functions which map from the current game state to (in this case) a boolean truth value, and those functions are then passed up the parse tree. Higher-level nodes, such as (and ...), receive the JAX functions corresponding to each of their children and return a new JAX function that also takes the game state as input and implements the appropriate operation (in this case, boolean conjunction). In pseudocode, using the Lark library’s Transformer paradigm, this looks like the following:
@@
In actuality, both the “children functions” and the combined “predicate function” must be written to be compatible with JAX’s vectorization scheme and just-in-time (JIT) compilation. This imposes a number of implementational constraints, most notably that the size and shape of all arrays must be fixed at compile time. This means, for instance, that the dimensions of the “legal action mask” (and, hence, the size of the action space in general) cannot change as the game progresses and values like the number of iterations in a loop or the positions of a lookup mask must be pre-specified. Crucially, however, values that are determined during parsing (such as the number of children for a given node, or the value of any arguments) can be safely passed into compiled JAX functions as static constants. This fact is what allows Ludax to dynamically create JAX functions that nonetheless obey the constraints of vectorization and JIT compilation. At the top of the parse tree, these composed JAX functions are ultimately used to define the behaviors that appear in the environment’s step function, such as applying the player’s action to the board and handling move effects.
@@
Precomputation: An important optimization for hardware accelerated environments is to express functions as batched matrix operations rather than iterative procedures. For instance, rather than checking for a line of pieces in Tic-Tac-Toe by starting at the position of the last move and scanning out in each direction (as Ludii’s implementation does), PGX hard-codes the set of board indices that correspond to each possible line of three in the game (i.e. [[0, 1, 2], [0, 3, 6], ...]) and performs a single multi-dimensional index into the board array. Ludax adopts and generalizes this approach: during parsing of line, for example, the line indices are computed with respect to the size and shape of the game board (i.e. rectangular, hexagonal, …) as well as the length and orientation of the desired line (i.e. diagonal, vertical, …). Again, because these values depend only on attributes that are determined during parsing, they can be passed into JAX functions as constants and provide a substantial speedup when available.
@@
Dynamic State Attributes: Different games require tracking different kinds of information about the current game state (such as scores or captured pieces). When Ludax compiles a game, it automatically extracts the attributes required to instantiate a game state and omits the others, thereby reducing the memory footprint of the entire state object. Ludax also automatically adds intermediary operations to each call of the environment’s step function that help speed up later computations. For example, Hex requires finding the connected components of the board to determine whether opposite edges are linked by a player’s pieces. Rather than re-compute the connected components from scratch after each move, Ludax again adopts an approach from PGX by detecting the presence of the connected function and maintaining a separate array of connected components that is locally updated after each move. This general procedure can be used to accommodate games with atypical or computationally expensive rules without affecting the runtime of other games.
@@
Figure 2: Average throughput (moves per second) on various exemplar games for Ludax, Ludii, and PGX. The first four games are implemented in all three frameworks, while the remaining games are implemented only in Ludax and Ludii. Speeds for Ludax and PGX are reported for 500 episodes of various batch sizes on a workstation with a single NVIDIA 4090 GPU and 36 threads on 18 CPU cores, while speeds for Ludii are the fasted recorded throughput for parallel execution on the same workstation across 1, 16, and 32 threads. Error bars are standard deviations over the 500 episodes.